OpenStudy (anonymous):

Complete the square: x2 + 2x + ____ . A. 1/2 B. 1 C. 2 D. 4

Parth (parthkohli):

I am not gonna go with that boring approach; let's try to be logical.

Parth (parthkohli):

OK, what is \((a + b)^2\)?

OpenStudy (anonymous):

1, x2+2.x.1+1^2= (x+1)^2

Parth (parthkohli):

@tanvirms You really shouldn't answer... please.

Parth (parthkohli):

Answering questions has a very bad impact.

OpenStudy (anonymous):

well i answered cuz he wanted an answer :)

Parth (parthkohli):

Dang, no! Don't ever give anyone the answer. Help them by leading them to an answer and explaining.

Parth (parthkohli):

@Luis_Rivera I take it that you meant\[x^2 + 2x + \left(2\over 2\right)^2 = x^2 + 2x + 1 \]